# P3Nav: A Unified Framework for Embodied Navigation Integrating Perception, Planning, and Prediction

### v1

@@ 3.1.3 LLM
Finally, we integrate the position-enhanced historical features $\mathbf{V_{p}}$ with the current observations $\mathbf{UV_{t}}$ and $\mathbf{X_{t}}$ from the visual encoder to create visual tokens for the LLM. In this framework, $\mathbf{UV_{t}}$ and $\mathbf{X_{t}}$ capture the agent’s ongoing environment, while $\mathbf{V_{p}}$ offers context from past observations, effectively filtering out redundant positions. Similarly, input instructions or questions are transformed into language tokens $\mathbf{L}\in\mathbb{R}^{n_{\text{L}}\times c}$ via a tokenizer, where $n_{\text{L}}$ denotes the length of the language tokens. These visual and language tokens are combined and fed into the LLM. The LLM utilizes its abilities in multimodal alignment and comprehension to process these tokens, decoding the outputs through specialized heads. For navigation datasets, it employs an action head to generate the executable action, while for EQA datasets, it uses an answer head (i.g., LLM head) to produce the natural language answer to input questions. Please refer to the supplementary material for more details.
@@ 3.2 Embodied Question Answering
Instruction-centric question. We construct an instruction-centric question that helps to generate comprehensive first-person analysis for navigation objectives (e.g., “go to a laptop in the bedroom”), through two synergistic mechanisms: role specification (“assuming you are a robot”) and exemplar-based prompting (i.e., two examples of navigation reasoning and planning, highlighted in red in Fig. 4). This dual approach ensures structured outputs that contain three essential components: (1) surrounding scene description (e.g., recognizing bedroom furniture layouts), (2) path planning analysis (e.g., creating obstacle-avoidance trajectories), and (3) common sense reasoning (e.g., prior knowledge about where laptops are typically placed).
Spatiotemporally coherent context construction. To ensure alignment with the adaptive 3D-aware history sampling strategy established in section 3.1.2, we systematically select the final $W$ frames from each navigation trajectory. This frame range captures critical target-approaching phases while maintaining temporal consistency with upstream visual processing modules. By adopting this temporal window, we construct spatiotemporally coherent contexts that encode successful navigation patterns and preserve motion dynamics essential for trajectory analysis.
Structured responses. Building upon the prepared inputs above, we synthesize instruction-aligned multimodal prompts by combining textual question with curated visual sequences. This integration requires GPT-4o [12] to generate structured responses through three distinct reasoning phases. First, the surrounding scene description identifies architectural features (e.g., “the room has a dark floor and walls”) and object semantics (e.g., “a yellow/blue bedspread”). Second, path planning analysis devises hierarchical navigation strategies (e.g., “enter bedroom → circumvent bed → final approach”). Third, common sense reasoning incorporates human-environment interaction norms (e.g., “laptops are commonly found on desks, dressers, or nightstands”). Each phase operates sequentially, ensuring explicit alignment between perception, action planning, and contextual knowledge.
@@ 3.3 Multitask Collaboration
Figure 5: Left: Both P3Nav and SPOC [10] generate the shortest path when the agent is near the target (e.g., in the same room). Right: In contrast, SPOC [10] fails due to repeatedly searching along the same paths when distant from the target (e.g., in different rooms), whereas P3Nav succeeds by effectively avoiding revisiting areas and exploring new ones.
In contrast, during the inference phase, the model focuses exclusively on predicting navigation actions. By maintaining this setting, the model effectively utilizes the enhanced perceptual insights and planning strategies gained during training, enabling efficient generation of precise navigation actions. Please refer to the supplementary material for more details of training objective.
@@ 4.1 Experimental Settings
Dataset and Metrics. We evaluate our proposed method on the $\mathrm{CHORES}$-$\mathbb{S}$ ObjectNav benchmark [10]. This benchmark includes 15 object categories and annotates 99k trajectories within 10k training houses, among 5M expert trajectory frames in the AI2-THOR simulated environment [14]. Besides, we extend the navigation dataset with EQA pairs for joint training, as described in Sec. 3.2. We follow the official evaluation split of the $\mathrm{CHORES}$-$\mathbb{S}$ ObjectNav benchmark [10], which contains 200 trajectories in 200 testing houses. To comprehensively demonstrate the effectiveness of our method, we evaluate performance using three metrics: success rate (SR), episode-length weighted success (SEL) [9], and percentage of rooms visited (%Rooms) [10]. The additional explanations of the metrics and implementation details are provided in the supplementary materials.
@@ 4.2 Performance Comparisons
As shown in Table 1, we compare our method with EmbSigLIP [10], SPOC [10], and $\text{SPOC}^{*}$ [10]. The performance results are directly copied from the original paper or supplementary materials, with “-” indicating that the results have not been disclosed. EmbSigLIP [10] denotes an implementation of the EmbCLIP [13] model that has been upgraded to utilize the SigLIP [31] backbone, serving as a baseline model. SPOC [10] focuses on imitating shortest paths for effective navigation by modeling optimal routes. $\text{SPOC}^{*}$ [10] is akin to SPOC [10], but it is trained using a larger set of expert trajectories. P3Nav integrates perception, planning, and prediction by jointly training on navigation and EQA tasks using the multitask collaboration strategy and employs the adaptive 3D-aware history sampling strategy to process historical frames.
@@
As presented in Table 1, EmbSigLIP [13] demonstrates limited navigational performance, with the lowest SR of 36.5 and a SEL of 24.5, highlighting inefficiencies in path length. SPOC [10] shows improvement, achieving an SR of 57.0 through optimal route modeling, and obtains the highest SEL of 46.2 by focusing on shortest path learning. $\text{SPOC}^{*}$ [10] further enhances SR to 60.0 via more expert trajectory training, although it results in longer paths, reducing SEL to 30.5. P3Nav utilizes adaptive 3D-aware history sampling and multitask collaboration strategies to achieve the highest SR of 75.0, effectively leveraging historical data for improved path decision-making and enhancing environmental understanding through EQA tasks. The results demonstrate that P3Nav excels in long-term navigation, surpassing other methods by integrating planning skills across both navigation and EQA tasks.
4.3 Qualitative Results
To further illustrate the effectiveness of our unified framework for embodied navigation, some visualization results compared with SPOC [10] are illustrated in Fig. 5. In the left case, both P3Nav and SPOC [10] are capable of generating the shortest path when the agent is positioned close to the target, such as being in the same room. Both methods demonstrate a strong understanding of the environment, allowing them to navigate efficiently without significant difficulties. However, the right case illustrates that when the agent is farther from the target, such as in different rooms, SPOC [10] faces challenges by repeatedly searching the same paths, leading to inefficiencies in navigating complex environments. In contrast, P3Nav avoids revisiting explored areas, enabling new path exploration and enhanced navigation efficiency, thus excelling in long-term navigation.
4.4 Ablation Studies
Does the adaptive 3D-aware history sampling help? We present an ablation study examining three key components, as shown in Table 2, which evaluates the impact of historical observations (HO), sampling (SP), and position-enhanced historical features (PEHF) in adaptive 3D-aware history sampling. Exp. 1, with no historical observations (i.e., window size $W=0$), focuses solely on the current visual information and achieves the lowest SR among the four experiments, highlighting the importance of historical observations for navigation. Comparing Exp. 2 and 4, we observe a 2.9 decrease in SR when position-enhanced historical features are not used, indicating that these features effectively avoid repeated exploration of the same area by adding positional information, thus enhancing navigation efficiency. Between Exp. 3 and 4, the absence of sampling in historical observations (i.e., threshold $\epsilon=0$) results in a 1.8 drop in SR, demonstrating that sampling helps improve navigation accuracy by effectively removing redundant observations.
@@ 4.4 Ablation Studies
Does the multitask collaboration strategy help? We conduct an ablation study on the multitask collaboration strategy to evaluate the benefits of joint training on both navigation and EQA tasks. As demonstrated in Table 3, Exp. 2, which additionally utilizes the EQA dataset during the training phase, achieves an increase of 3.0 in SR and 4.0 in SEL compared to Exp. 1, which only employs the navigation dataset. This result indicates that integrating the EQA dataset allows agents to leverage perception and planning to improve their prediction capabilities, thereby enhancing navigation abilities. Moreover, we have identified a remarkable improvement: Exp. 4, trained on the ObjectNav benchmark, which contains five times more samples than ObjectNavRoom, exhibits a 9.3 increase in SR compared to Exp. 3. This suggests that increasing the volume of EQA data can significantly enhance navigation performance.
@@
1 ObjectNavRoom 42.1 15.2 75.0
2 ObjectNavRoom 45.1 19.2 71.9
3 ObjectNav 65.7 31.2 68.4
4 ObjectNav 75.0 34.3 69.4
@@
On hyper-parameters. We investigate the performance using different values of the window size $W$ and the threshold $\epsilon$ for the adaptive 3D-aware history sampling strategy, as they control the length of historical context and the sampling density, respectively. Through a coarse search shown in Table 4, we adopt $W=60$ and $\epsilon=0.1$ for our adaptive 3D-aware history sampling strategy.
Exp $W$ $\epsilon$ SR $\uparrow$ SEL $\uparrow$ %Rooms
1 20 0.1 32.6 13.2 74.8
2 40 0.1 34.8 13.0 74.7
3 60 0.1 42.1 15.2 75.0
4 80 0.1 38.9 16.7 74.3
5 100 0.1 35.2 14.3 72.3
6 60 0.05 41.5 10.3 79.9
7 60 0.15 41.3 12.6 77.8
8 60 0.2 40.5 13.7 72.1
@@ 5 Conclusion
In this paper, we present the P3Nav framework, which significantly advances language-guided visual navigation by effectively integrating Perception, Planning, and Prediction through Multitask Collaboration strategy. Additionally, P3Nav adopts an Adaptive 3D-aware History Sampling strategy to enhance navigation by efficiently utilizing historical data to minimize redundancy. The framework’s impressive 75% success rate in ObjectNav on the $\mathrm{CHORES}$-$\mathbb{S}$ benchmark underscores its effectiveness. Although P3Nav has achieved remarkable results in simulated environments, there is still a long way to go to achieve truly human-like embodied navigation. Future work will address these limitations by exploring the integration of larger, more diverse datasets to further enhance P3Nav’s generalization capabilities across various environments.
@@ 6 Implementation Details
In this section, we elaborate on the more detailed implementations for P3Nav in Sec. 3.
@@ 6.1.1 Vison Encoder
Following the settings in previous works [17, 28], we input RGB images from both the head perspective $I_{t}^{head}$ and the wrist perspective $I_{t}^{wrist}$ into ViT [19] to obtain the 2D features $X_{t}^{head}$ and $X_{t}^{wrist}$. Both $X_{t}^{head}$ and $X_{t}^{wrist}$ are then fed into UVFormer [17] to construct multi-view 3D features $UV_{t}$. In the navigation task, we adhere to the settings in [17, 28], utilizing only the wrist perspective $X_{t}^{wrist}$ as the 2D feature $X_{t}$ to broaden the exploration view. Conversely, in the EQA task, we use the head perspective $X_{t}^{head}$ for the 2D feature $X_{t}$ to maintain a first-person perspective, as the EQA pairs are generated from this perspective.
@@ 6.1.2 LLM
We utilize OpenFlamingo [2] as our LLM, freezing the self-attention layers during training while fine-tuning the cross-attention layers. For the action head, we employ a multi-layer perceptron to map the final hidden states produced by the LLM from the $c$-dimensional space to the action space of the $\mathrm{CHORES}$-$\mathbb{S}$ ObjectNav benchmark [10]. For the answer head (i.e., LLM head), we apply the argmax operation on the logits output by the LLM to decode the answer.
@@ 6.2 Training Objective
Action prediction Loss. We utilize behavior cloning to train the navigation model. Given an expert trajectory $\tau=\left(\hat{a_{0}},\cdots,\hat{a_{T}}\right)$, we use the cross-entropy loss for action prediction. The loss for the trajectory is as follows:

### v4

@@ 3.1.3 LLM
Finally, we integrate the position-enhanced historical features $V_{p}$} with the current observations $UV_{t}$} and $X_{t}$} from the visual encoder to create visual tokens for the LLM. In this framework, $UV_{t}$} and $X_{t}$} capture the agent’s ongoing environment, while $V_{p}$} offers context from past observations, effectively filtering out redundant positions. Similarly, input instructions or questions are transformed into language tokens ${L}\in\mathbb{R}^{n_{\text{L}}\times c}$ via a tokenizer, where $n_{\text{L}}$ denotes the length of the language tokens. These visual and language tokens are combined and fed into the LLM. The LLM utilizes its abilities in multimodal alignment and comprehension to process these tokens, decoding the outputs through specialized heads. For navigation datasets, it employs an action head to generate the executable action, while for EQA datasets, it uses an answer head (i.g., LLM head) to produce the natural language answer to input questions.
@@ 3.2 Embodied Question Answering
Instruction-centric question. We construct an instruction-centric question to elicit comprehensive first-person navigation analysis (e.g., “go to a laptop in the bedroom”) using two synergistic mechanisms: role specification (“assuming you are a robot”) and exemplar-based prompting (two navigation reasoning and planning examples, see Fig. 4). This ensures structured outputs covering: (1) scene description (e.g., bedroom furniture layout), (2) path planning (e.g., obstacle avoidance), and (3) common sense reasoning (e.g., typical laptop locations).
Spatiotemporally coherent context construction. To ensure alignment with the adaptive 3D-aware history sampling strategy established in Sec. 3.1.2, we systematically select the final $W$ frames from each navigation trajectory. This frame range captures critical target-approaching phases while maintaining temporal consistency with upstream visual processing modules. By adopting this temporal window, we construct spatiotemporally coherent contexts that encode successful navigation patterns and preserve motion dynamics essential for trajectory analysis.
Structured responses. Building on the prepared inputs, we synthesize instruction-aligned multimodal prompts by combining textual questions with curated visual sequences. GPT-4o [14] then generates structured responses in three phases. First, the surrounding scene description identifies architectural features (e.g., “the room has a dark floor and walls”) and object semantics (e.g., “a yellow/blue bedspread”). Second, path planning analysis devises hierarchical navigation strategies (e.g., “enter bedroom → circumvent bed → final approach”). Third, common sense reasoning incorporates human-environment interaction norms (e.g., “laptops are commonly found on desks, dressers, or nightstands”). Each phase operates sequentially, ensuring explicit alignment between perception, action planning, and contextual knowledge.
@@ 3.3 Multitask Collaboration
During inference, the model focuses solely on predicting navigation actions, effectively leveraging the perceptual and planning abilities acquired during training for precise and efficient navigation. Please refer to the supplementary material for details on the training objective.
Figure 5: Left: Both RoboTron-Nav and SPOC [10] generate the shortest path when the agent is near the target (e.g., in the same room). Right: In contrast, SPOC [10] fails due to repeatedly searching along the same paths when distant from the target (e.g., in different rooms), whereas RoboTron-Nav succeeds by effectively avoiding revisiting areas and exploring new ones.
@@ 4.1 Experimental Settings
Dataset and Metrics.
We evaluate our proposed method on the $\mathrm{CHORES}$-$\mathbb{S}$ ObjectNav benchmark [10]. Besides, we extend the navigation dataset with EQA pairs for joint training, as described in Sec. 3.2. We follow the official evaluation split of the $\mathrm{CHORES}$-$\mathbb{S}$ ObjectNav benchmark [10], which contains 200 trajectories in 200 testing houses. To comprehensively demonstrate the effectiveness of our method, we evaluate performance using three metrics: success rate (SR), episode-length weighted success (SEL) [8], and percentage of rooms visited (%Rooms) [10]. The additional explanations of the metrics and implementation details are provided in the supplementary materials.
@@ 4.2 Performance Comparisons
As shown in Table 1, we compare our method with PIRLNav [23], JSRL [28], EmbSigLIP [10], SPOC [10], $\text{SPOC}^{*}$ [10], and RING [9]. PIRLNav [23] and JSRL [28] are first pretrained with imitation learning, and then fine-tuned via reinforcement learning with sparse rewards. EmbSigLIP [10] denotes an EmbCLIP [15] model upgraded with a SigLIP [35] backbone. SPOC [10] imitates shortest paths by modeling optimal routes, while $\text{SPOC}^{*}$ [10] is trained on a larger set of expert trajectories. RING [9] introduces a universal indoor navigation policy transferable across diverse robots without additional fine-tuning.
@@
As presented in Table 1, PIRLNav [23] and JSRL [28] achieve relatively low SRs (20.0 and 21.0), which can be attributed to the challenges of fine-tuning with sparse reward signals in reinforcement learning. EmbSigLIP [15] also demonstrates limited navigation performance (SR 36.5, SEL 24.5). SPOC [10] shows improvement, achieving an SR of 57.0 and a SEL of 46.2 by focusing on shortest path learning. $\text{SPOC}^{*}$ [10] further increases SR to 60.0 via additional expert trajectory training, though at the cost of longer paths, reducing SEL to 30.5. RING [9] achieves a much higher SR of 72.1 and the best SEL of 53.0, demonstrating strong generalization and efficiency. RoboTron-Nav leverages adaptive 3D-aware history sampling and multitask collaboration to achieve the highest SR of 81.1, effectively utilizing historical data for improved path decision-making and enhancing environmental understanding through EQA tasks. These results demonstrate that RoboTron-Nav excels in long-term navigation, surpassing other methods by integrating planning across both navigation and EQA tasks.
@@
1 ObjectNavRoom 42.5 16.0 62.3
2 ObjectNavRoom 47.2 21.9 65.8
3 ObjectNav 72.5 33.5 68.4
4 ObjectNav 81.1 43.9 67.8
@@
4.3 Qualitative Results
To further illustrate the effectiveness of our unified framework for embodied navigation, some visualization results compared with SPOC [10] are illustrated in Fig. 5. In the left case, both RoboTron-Nav and SPOC [10] are capable of generating the shortest path when the agent is positioned close to the target, such as being in the same room. Both methods demonstrate a strong understanding of the environment, allowing them to navigate efficiently without significant difficulties. However, the right case illustrates that when the agent is farther from the target, such as in different rooms, SPOC [10] faces challenges by repeatedly searching the same paths, leading to inefficiencies in navigating complex environments. In contrast, RoboTron-Nav avoids revisiting explored areas, enabling new path exploration and enhanced navigation efficiency, thus excelling in long-term navigation.
Exp $W$ $\epsilon$ $\tau$ SR $\uparrow$ SEL $\uparrow$ %Rooms
1 20 0.1 0 32.6 13.2 74.8
2 40 0.1 0 34.8 13.0 74.7
3 60 0.1 0 42.1 15.2 75.0
4 80 0.1 0 38.9 16.7 74.3
5 100 0.1 0 35.2 14.3 72.3
6 60 0.05 0 41.5 10.3 79.9
7 60 0.1 0 42.1 15.2 75.0
8 60 0.15 0 41.3 12.6 77.8
9 60 0.2 0 40.5 13.7 72.1
10 60 0.1 0.9 42.3 15.6 64.2
11 60 0.1 0.95 42.5 16.0 62.3
12 60 0.1 0.99 41.5 14.6 73.2
@@ 4.3 Qualitative Results
4.4 Ablation Studies
Does the adaptive 3D-aware history sampling help? As shown in Table 2, we conduct an ablation study on three key components in adaptive 3D-aware history sampling: historical observations (HO), sampling (SP), and position-enhanced historical features (PEHF). Exp. 1, which uses no historical observations ($W=0$), relies solely on current visual input and yields the lowest SR, underscoring the value of historical information. Comparing Exp. 2 and 4, removing PEHF leads to a 2.9 decrease in SR, indicating that positional information helps avoid redundant exploration and improves efficiency. Similarly, between Exp. 3 and 4, omitting sampling in historical observations (i.e., $\epsilon=0$, $\tau=1$) causes a 2.2 drop in SR, showing that sampling reduces redundancy and enhances navigation accuracy.
Does the multitask collaboration strategy help? We conduct an ablation study on the multitask collaboration strategy to evaluate the benefits of joint training on both navigation and EQA tasks. As demonstrated in Table 3, Exp. 2, which additionally utilizes the EQA dataset during the training phase, achieves an increase of 4.7 in SR and 5.9 in SEL compared to Exp. 1, which only employs the navigation dataset. This result indicates that integrating the EQA dataset allows agents to leverage perception and planning to improve their prediction capabilities, thereby enhancing navigation abilities. Moreover, we have identified a remarkable improvement: Exp. 4, trained on the ObjectNav benchmark, which contains five times more samples than ObjectNavRoom, exhibits a 8.6 increase in SR compared to Exp. 3. This suggests that increasing the volume of EQA data can significantly enhance navigation performance.
Effect of hyper-parameters. We investigate the performances using different values of the window size $W$, the relative position threshold $\epsilon$, and the semantic similarity threshold $\tau$ for the adaptive 3D-aware history sampling strategy. Here, $W$ controls the length of historical context, while $\epsilon$ and $\tau$ jointly determine the density and diversity of historical sampling in the spatial and semantic domains, respectively. As shown in Table 4, we conducted comprehensive grid search experiments to analyze specifically how the window size $W$ (20-100), relative position threshold $\epsilon$ (0.05-0.2), and semantic similarity threshold $\tau$ (0.9-0.99) affect performance. Increasing $W$ initially improves performance by providing richer historical information, with the best results at $W=60$ (around 6 meters). However, when $W$ exceeds 60, the inclusion of excessive history starts to overwhelm the model, leading to reduced performance. For spatial sampling, increasing $\epsilon$ up to 0.1 improves performance by reducing spatial redundancy through lowering the sampling frequency, with optimal results at $\epsilon=0.1$. However, further increases make sampling overly sparse and reduce performance. For semantic sampling, increasing $\tau$ up to 0.95 improves performance by preserving useful contextual information, with optimal results at $\tau=0.95$. However, further increases introduce excessive redundancy and reduce performance. Therefore, we adopt $W=60$, $\epsilon=0.1$, and $\tau=0.95$ for our adaptive 3D-aware history sampling strategy.
@@ 5 Conclusion
In this paper, we present the RoboTron-Nav framework, which significantly advances language-guided visual navigation by effectively integrating perception, planning, and prediction through multitask collaboration strategy. Additionally, RoboTron-Nav adopts an adaptive 3D-aware history sampling strategy to enhance navigation by efficiently utilizing historical data to minimize redundancy. The framework’s impressive 81.1% success rate in ObjectNav on the $\mathrm{CHORES}$-$\mathbb{S}$ benchmark underscores its effectiveness. Although RoboTron-Nav has achieved remarkable results in simulated environments, there is still a long way to go to achieve truly human-like embodied navigation. Future work will address these limitations by exploring the integration of larger, more diverse datasets to further enhance RoboTron-Nav’s generalization capabilities across various environments.
@@ 6 Implementation Details
In this section, we elaborate on the more detailed implementations for RoboTron-Nav in Sec. 3.
@@ 6.1.1 Vison Encoder
Following the settings in previous works [19, 32], we input RGB images from both the head perspective $I_{t}^{head}$ and the wrist perspective $I_{t}^{wrist}$ into ViT [22] to obtain the 2D features $X_{t}^{head}$ and $X_{t}^{wrist}$. Both $X_{t}^{head}$ and $X_{t}^{wrist}$ are then fed into UVFormer [19] to construct multi-view 3D features $UV_{t}$.
We adopt UVFormer [19] as our 3D occupancy predictor. As shown in Eq. (2), UVFormer takes the image features $X_{t}$, camera parameters $Cam$, and a set of learnable UniView queries $Q$ as input, and outputs a unified view representation $UV_{t}$. The query set $Q=\{Pos,Emb\}$ comprises positional encodings $Pos\in\mathbb{R}^{L\times B\times 3P}$ and learnable embeddings $Emb\in\mathbb{R}^{L\times B\times C}$. Here, $L$ and $B$ (both set to $20$) specify the 3D grid’s spatial layout within the robot’s workspace, and $P$ is the number of uniformly sampled points along the vertical axis of each pillar cell. Each pillar cell covers $0.05^{2}$ square meters on the ground and spans $0.5$ meters in height. $Emb^{l,b}\in\mathbb{R}^{C}$ encodes features for each pillar cell. The camera parameters $Cam$ correspond to $N$ different viewpoints. The unified view representation $UV_{t}\in\mathbb{R}^{L\times B\times C}$ integrates information from the entire $L\times B\times P$ 3D grid and serves as the basis for downstream occupancy prediction.
In the navigation task, we adhere to the settings in [19, 32], utilizing only the wrist perspective $X_{t}^{wrist}$ as the 2D feature $X_{t}$ to broaden the exploration view. Conversely, in the EQA task, we use the head perspective $X_{t}^{head}$ for the 2D feature $X_{t}$ to maintain a first-person perspective, as the EQA pairs are generated from this perspective.
@@ 6.1.2 LLM
We utilize MPT [^1] https://huggingface.co/mosaicml/mpt-1b-redpajama-200b-dolly as our LLM, freezing the self-attention layers during training while fine-tuning the cross-attention layers. For the action head, we employ a multi-layer perceptron to map the final hidden states produced by the LLM from the $c$-dimensional space to the action space of the $\mathrm{CHORES}$-$\mathbb{S}$ ObjectNav benchmark [10]. For the answer head (i.e., LLM head), we apply the argmax operation on the logits output by the LLM to decode the answer.
@@ 6.2 Training Objective
[…]